# Portable Biomechanics Laboratory: Clinically Accessible Movement Analysis from a Handheld Smartphone

### v1

@@ Results
For the MMMC dataset, the MJAE between a handheld moving camera and MMMC was 2.79°. This error was slightly higher for static camera fits using a single view from the MMMC system (2.96°), likely due to the greater distances to the person. The error was the greatest in the neurological population, with an MJAE of 3.32 degrees, lower for LLPUs (2.97°), and lowest for controls at 2.51 degrees (Fig. 4B, Table S2). Our method also performed well on the BML-MoVi dataset, showing only 2.74 degrees of MJAE from the monocular fits compared to OMC.
Frontal-plane joint angles (e.g., hip adduction) were best estimated from frontal views, while sagittal-plane joint angles (e.g., hip and knee flexion) were more accurately estimated from sagittal views (one-sided Mann-Whitney U-tests; p$<$0.005 for all comparisons). This difference was greatest for knee flexion at 1.02 degrees (Fig. 3C).
Visual comparisons of root position between single-camera and multi-view fits are shown in Figure 3D. The clinical dataset using a handheld, rotating smartphone yielded an average RTE of 6.44 cm, while a static camera yielded 5.00 cm. RTE for the BML-MoVi dataset with a static smartphone averaged 2.41 cm (Table 1).
Figure 3: Quality Measures of Single Camera Fitting A) Kinematic traces from smartphone video (red/blue) compared to ground truth (gray dashed) during walking. B) Joint angle errors across populations for select lower limb angles. n denotes the number of unique individuals in each cohort and v denotes the number of total videos for that cohort. C) Select joint angle errors with respect to camera view angle show that sagittal plane angles have the lowest error with sagittal camera views, and frontal angles have the lowest error with frontal views. D) Pelvis translation (RTE) extracted from handheld smartphone video compared to ground truth during functional gait assessments.
Figure 4: Joint Angle Error Distributions. A) In the MMMC dataset, MJAE distributions remain relatively stable when aggregated across activity types, participant populations (B), as well as between handheld (moving) and static cameras. C) Joint-specific MJAE distributions from both the MMMC and BML-MoVi datasets.
Table 1: Summary of errors for different datasets. MJAE is the median joint angle error (in degrees) and RTE is the median root translation error (in cm). Errors are presented as: (Median (nIQR)).
@@
MJAE (deg) 2.79 (0.86) 2.96 (0.78) 2.74 (0.50)
@@
Movement Quality Metrics
The Gait Deviation Index (GDI) provides a holistic measure of movement quality that we use to summarize the information found in PBL-based measures of movement. We test this measure for repeatability, validity, and responsiveness.
GDI Repeatability
We analyzed the $>$13,500 steps across 222 individuals recorded with the PBL system, and found high repeatability of 0.84 for the GDI of a single step (ICC2) and higher repeatability when averaging multiple steps together: 0.96 (ICC2k) (Table 2). ICCs for spatiotemporal metrics were lower—likely due to timing variability—but remained highly repeatable when averaged (ICC2k).
GDI Construct Validity
We found the GDI measured with our PBL is sensitive to clinical differences between participants (Fig. 5). Performance on the BBS is used to classify fall risk and was measured for our neurologic participants. The GDI was statistically different between the low and high fall-risk participants, and both were different than the control participants (Fig. 5B). The GDI also correlates well with walking speed measured during the 10MWT (Pearson’s $r=0.82$), one of the most commonly measured clinical outcomes for gait (Fig. 5C). It is important to note GDI and 10MWT measure different things, as it is possible to walk faster using a higher cadence with an unchanging kinematic trajectory. Next, the GDI is significantly decreased in transtibial amputees compared to controls and further decreased for transfemoral amputees (Fig. 5D). The KOA population was also statistically different from control with a mean GDI of 83.5 (10.8). Finally, the GDI showed a significant correlation with mJOA ($r=0.47$) and the lower extremity subscore of the mJOA ($r=0.48$) (Fig. 5E, Table 2).
GDI Responsiveness
We found that quantitative, PBL-based metrics all had higher responsiveness to surgical intervention than the mJOA, a commonly used clinical outcome measure (Table 2). Notably, the jackknife confidence intervals for GDI and DST did not cross zero while the mJOA did, suggesting greater effect size stability, though not necessarily statistical significance.
@@
Table 2: Reliability, Validity, and Responsiveness of Smartphone-Based Metrics. ICCs are presented as ICC (95% confidence intervals). mJOA and mJOA-LE show Spearman correlations between gait metric and ordinal ratings. SRM a unitless measure of change in metric before and after surgery, with 95% confidence intervals. * indicates $p<0.05$.
ICC2 ICC2k r mJOA r mJOA-LE SRM
mJOA - - - - 0.20 (-0.38 0.78)
mJOA-LE - - - - 0.14 (-0.45 0.72)
Cadence 0.58 (0.52 0.64) 0.88 (0.85 0.90) 0.25* 0.39* 0.34 (-0.20 0.88)
DST 0.41 (0.35 0.48) 0.78 (0.72 0.82) -0.27* -0.37* -0.48 (-0.88 -0.07)
GDI 0.84 (0.81 0.86) 0.96 (0.96 0.97) 0.47* 0.48* 0.45 (0.00 0.91)
Discussion
@@
We present a clinically validated method for extracting accurate biomechanics from smartphone video in clinical settings. Across diverse populations, our single-camera system showed strong agreement with ground truth (typically $<3^{\circ}$ error). Gait quality metrics derived from these videos detected clinically relevant group differences and were more responsive than the mJOA following decompression surgery. These findings suggest that a clinician, therapist, or medical assistant could feasibly perform gait analysis in under a minute to obtain an objective measurement of a patient’s gait impairment to sensitively track their recovery.
@@
Joint angle errors were generally low across all populations and settings (Table 1). We attribute the slightly higher MJAE in clinical populations (0.46 - 0.81 degrees, Fig. 3B, Fig. 4B, Table S2) to assistive devices such as canes, walkers, ankle foot orthoses, or gait belts (Fig. 2) and the fact that keypoint detection does not perform as well on prosthetic devices (30).
A Responsive and Valid Metric of Functional Status
Our method measured reliable, valid, and responsive features of movement which is arguably more important for clinical practice than joint angle accuracy. GDI demonstrated construct validity for multiple clinical populations, such as being different for inpatients with a high fall risk than low fall risk based on the BBS (Fig. 5B) and correlated highly with the 10MWT (Fig. 5C), both important outcome measures used commonly in clinical practice.
The GDI had a moderate correlation with the mJOA. We note that we do not necessarily expect a strong association between the GDI, a continuous quantitative measure, as the mJOA is a patient-reported ordinal scale. Despite the differences, the GDI and mJOA both responded positively to spinal decompression in the surgical group recorded before/after their surgery (Table 2). Notably, the GDI had a higher SRM (0.45) than both the mJOA (0.20) and mJOA-LE (0.14), with a 95% confidence interval that remained entirely above zero. This indicates that our easy-to-use system could be more sensitive than the widely used mJOA.
@@ A Responsive and Valid Metric of Functional Status
We focused on the GDI because it is a widely used summary of gait kinematics. However, we do not claim it is the optimal metric for every application. Instead, we anticipate that scalable gait analysis will enable new research to identify the most clinically meaningful measures. For example, we recently described a method for inferring torques and ground reaction forces from kinematics and saw that the ground reaction forces are sensitive to clinical history (48).
Limitations
@@
The minimally clinically important difference for kinematic features of gait has not yet been established for many clinical conditions. While our accuracy is generally quite good, greater accuracy may be required for certain indications. For example, the 5 degrees of error we typically see at the ankle might be insufficient to detect subtle changes in plantarflexion spasticity or toe clearance from a single camera.
We also note the heavy tails in our error distributions with some trials having errors $>$10 degrees (Fig. 4C) and our analysis of the reconstruction errors (Fig. S2) indicated little relationship between reconstruction error and MJAE, highlighting the need for reliable confidence scores from markerless motion capture. We have developed methods for both monocular videos without biomechanics (49), and for biomechanics from multi-view video (25). Extending these confidence intervals to monocular biomechanics is an important future direction.
Uniquely Accessible and Validated Approach for Clinical Implementation
Other approaches have proposed deployable systems for human motion capture. Most well-known is OpenCap, which uses two smartphones on static tripods (24) to triangulate joint locations from video and fit biomechanics using the widely used OpenSim framework (50, 51). During walking, the approach presented here shows similar joint angle errors to OpenCap (4.1°) using just a single, moving, camera. OpenCap does go a step further than our approach and produces joint kinetics, which we can incorporate in the future, such as using our Kinetic Twin approach (48). Our experience with PBL shows that the ability to use a single portable camera is critical for easy use in the clinic.
Two recent works estimate biomechanics from monocular video: BioPose (52) and Human Skeleton and Mesh Recovery (HSMR) (53). Neither was evaluated on clinical populations. Our method differs by using trajectory optimization with a single, fixed skeleton scale across trials, and it can readily incorporate wearable sensors (32). HSMR directly regresses joint angles from images, which is fast but likely limited to poses seen in its training data of able-bodied individuals—raising concerns about generalization to clinical movements. It also does not report joint angle errors, making comparisons difficult. On the BML-MoVi dataset, our method performs comparably to BioPose. Ultimately, a proliferation of AI-powered methods for accessible biomechanics will increase use of clinical movement analysis. These large datasets will enable data-driven precision rehabilitation interventions (54).
@@
In conclusion, we present a method for extracting accurate and interpretable joint kinematics from smartphone videos. We validate this method in inpatient, outpatient therapy, and clinical settings, finding it to be highly usable and robust—even in the presence of nearby therapists and assistive devices. Importantly, gait scores extracted from this approach correlate well with commonly used clinical scales such as the 10MWT, BBS, and mJOA. This approach enables increased fidelity in clinical and at-home monitoring of movement impairments, paving the way for a big data revolution in movement science.
@@ Acknowledgements.
This study was approved by the Northwestern University Institutional Review Board. Data were collected in our laboratory using handheld smartphone video (30), an instrumented GaitRite walkway, and an MMMC system (41). The dataset includes recordings from 33 control participants, 48 lower limb prosthesis users (LLPUs), and 73 individuals with a history of neurological injury, including 6 pediatric cases. All recording sessions consisted of overground walking along the GaitRite walkway (8 m) while a researcher followed the participant with smartphone.
@@
Portable Biomechanics Laboratory (PBL) All smartphone RGB video was captured on a Samsung Galaxy S8 or Samsung Galaxy S20 Ultra at 30 Hz and 1080 $\times$ 1920 resolution (30) using a custom application. This application simultaneously logs smartphone RGB video, and if available, depth video, internal phone sensor logs (gyroscope, magnetometer, accelerometer, and fused orientation estimate), and wearable sensor data (55). During walking trials, a researcher held the phone in portrait orientation mounted on a 3-axis gimbal and followed the participant from behind at a mostly sagittal, but partially oblique view. The gimbal was used solely to improve visualizations; the method remains accurate without it, though the resulting video may be slightly shakier. FSST was collected at a frontal view in front of the participant, and TUG was collected from a sagittal view with the camera rotating to follow the participant.
The video and sensor data from the smartphone application was uploaded to a secure cloud server for storage and later downloaded into PosePipe (35), an open source application implementing a variety of computer vision tasks such as person detection, annotation, and 2D and 3D keypoint detection. PosePipe tracks all data and analyses in a MySQL database using DataJoint (56). This makes it feasible to manage the 100s of thousands of videos we have collected.
Multi-View Markerless Motion Capture (MMMC) Recordings in the laboratory dataset employed the Multi-View Markerless Motion Capture (MMMC) system which consists of 8-12 FLIR BlackFly S GigE cameras which acquire synchronized RGB video at 29 fps. Cameras were arranged such that at least three cameras covered participants at all points during a recording. A detailed description of this system can be found in (41). For this system, we reconstructed biomechanical fits using methods described in (26). In total, the dataset includes 1,267,586 frames across 2,206 trials from 143 participants—comprising 1,056,861 frames from 2,028 trials of 116 participants recorded with both the PBL and MMMC systems, and an additional 210,725 frames from 178 trials of 27 participants recorded with a single MMMC camera—amounting to 11.7 hours of video.
OMC Dataset: BML-MoVi
The publicly available BML-MoVi (38) dataset contains synchronous optical motion capture (OMC) data and smartphone RGB video data of 90 able-bodied participants performing a variety of everyday and sports movements. The OMC data contains a dense set of 87 markers, suitable for reconstructing biomechanics recorded at 120 Hz.
For the OMC data, we fit our model using only the 3D keypoint loss as the markers had already been triangulated and cleaned, following the approach from (26). For the smartphone data, we apply the 3D and 2D keypoint loss following the method we use on the PBL data and describe below.
@@
This study was approved by the Northwestern University and Washington University in St. Louis Institutional Review Boards. To validate our methods in a real-world clinical environment, we deployed our PBL system (30) in two outpatient clinics. The first was a neurosurgery clinic treating patients with cervical myelopathy (CM)—a degenerative condition of the cervical spine that often impairs gait and balance. During routine pre- and post-operative visits for spinal decompression, participants were asked to walk three times at a self-selected pace and three times at a fast pace down the clinic hallway while being recorded from behind by a clinician. At each visit, patients completed the modified Japanese Orthopedic Association (mJOA) questionnaire (57, 58) to assess their symptoms. The CM cohort includes 911 videos from 55 patients collected at multiple timepoints: pre-surgery, 6 weeks post-surgery, 3 months post-surgery, and 1 year post-surgery. 17 of these participants have sessions before and after surgery, which were used in our sensitivity analysis.
@@
We employed MeTRAbs-ACAE (37) to detect 2D and 3D virtual marker locations from RGB videos. As recommended by the author, we approximated confidence by measuring the standard deviation of each 3D joint location estimated from 10 different augmented versions of each video frame. This was converted to a confidence estimate using a sigmoid function with a half maximum at 30 mm and a width of 10 mm. This results in a confidence score $c(t,j)\in[0,1]$ for each timepoint $t$ and joint $j$. For training, we only used frames where the person was fully in the video view, allowing for short periods ($<$1.0) seconds of partial coverage, reducing total dataset size to 12.4 hours.
@@
Smartphone and MMMC reconstructions utilized a biomechanically-grounded model implemented in Mujoco from LocoMujoco (59), originally based on an OpenSim model (60). Mujoco’s GPU acceleration engine, MJX, supports parallelizing forward kinematic passes of this model. We have previously optimized site locations for the 87 MoVi keypoints on a similar model (39). We also removed collisions other than the feet, added a neck joint with 3 degrees of freedom, and extended the hip flexion and extension ranges. Model scaling is controlled by 8 scaling parameters: overall size, the pelvis, left thigh, left leg and foot, right thigh, right leg and foot, the left arm, and the left leg. We follow notation used by the skinned multi-person linear model (SMPL) (61) and represent our GPU-accelerated, biomechanically-ground, forward kinematic equation as:
@@
We represent the kinematic trajectory for each trial as a learned implicit function, $f_{\theta}$, implemented as a multi-layer perceptron that takes time as an input using sinusoidal positional encoding (62) and outputs joint angles $\theta(t)$. In addition to the 40 joint angles output from the final layer, we also output 3 parameters representing the orientation of the smartphone in global space $\mathbf{r}$.

### v2

@@ Results
For the MMMC dataset, the Median Absolute Error between a handheld moving camera and MMMC was 2.79°. This error was slightly higher for static camera fits using a single view from the MMMC system (2.96°), likely due to the greater distances to the person. The error was the greatest in the neurological population, with a Median Absolute Error of 3.32°, lower for LLPUs (2.97°), and lowest for controls at 2.51° (Fig. S1, Table S2). Our method also performed well on the BML-MoVi dataset, showing only 2.74° of Median Absolute Error from the monocular fits compared to OMC. Error distributions are shown for lower limb joints in Figure 4B and broken down across all joints in Tables S1, S2 and Figure S1. Notably wide error distributions are Neck Extension, Lumbar Extension, Hip Flexion (static camera), with third quartiles above 5°.
Frontal-plane joint angles (e.g., hip adduction) were best estimated from frontal views, while sagittal-plane joint angles (e.g., hip and knee flexion) were more accurately estimated from sagittal views (one-sided Mann-Whitney U-tests; p<<0.005 for all comparisons). This difference was greatest for knee flexion at 1.02° (Fig. 4C).
Visual comparisons of root position between single-camera and multi-view fits are shown in Figure S3. The clinical dataset using a handheld, rotating smartphone yielded an average RTE of 6.44 cm, while a static camera yielded 5.00 cm. RTE for the BML-MoVi dataset with a static smartphone averaged 2.41 cm (Table 1).
Table 1: Summary of errors for different datasets. Median and Mean Absolute Errors are for joint angles aggregated across all joints, participants, and trials. RTE is the median root translation error (in cm). Errors are presented as: (Median (nIQR)).
@@
Median Absolute Error (deg) 2.79 (0.86) 2.96 (0.78) 2.74 (0.50)
Mean Absolute Error (deg) 4.10 (3.71) 4.49 (1.99) 3.89 (1.66)
@@
Figure 4: Quality Measures of Single Camera Fitting. A) Kinematic traces from smartphone video (red/blue) compared to ground truth (gray dashed) during walking. B) Joint angle errors across populations for select lower limb angles. Control: n=29 participants, v=950 videos, Neurologic: n=36, v=290, LLPU: n=40, v=458, Pediatric: n=4, v=25. C) Select joint angle errors with respect to camera view angle show that sagittal plane angles have the lowest error with sagittal camera views, and frontal angles have the lowest error with frontal views.
@@
Table 2: Reliability, Validity, and Responsiveness of Smartphone-Based Metrics. ICCs are presented as ICC (95% confidence intervals). mJOA and mJOA-LE show Spearman correlations between gait metric and ordinal ratings. SRM a unitless measure of change in metric before and after surgery, with 95% confidence intervals. Wilcoxon Effect Size presented with 95% confidence intervals. * indicates $p<0.05$.
Repeatability Validity Responsiveness
ICC2 ICC2k r mJOA r mJOA-LE Effect Size
Patient-Reported mJOA - - - - 0.24 (-0.35 0.76)
mJOA-LE - - - - 0.22 (-0.41 0.91)
Smartphone-Based Cadence 0.58 (0.52 0.64) 0.88 (0.85 0.90) 0.25* 0.39* 0.39 (-0.09 0.77)
DST 0.41 (0.35 0.48) 0.78 (0.72 0.82) -0.27* -0.37* -0.52 (-0.84 -0.11)*
GDI 0.84 (0.81 0.86) 0.96 (0.96 0.97) 0.47* 0.48* 0.44 (-0.02 0.76)*
Clinical Inference from Smartphone-Based Kinematics
Extracting valid, repeatable, and responsive measurements from our smartphone-based system is arguably more important than accuracy alone. We computed the GDI as a measure of gait quality, testing it for repeatability, validity and responsiveness (Figure 5, Table 2). We also test for differences in raw kinematic traces pre/post surgery using the paired SPM1D t-tests [pataky_generalized_2010, pataky_one-dimensional_2012].
Gait Measurement Clinical Validity
GDI separated groups with expected differences, for instance GDI was significantly lower for participants scoring higher fall risk on the Berg Balance Scale and Transfemoral LLPUs had a significantly lower GDI than Transtibial LLPUs (Fig. 5B,D). The GDI also correlates well with the 10MWT (Pearson’s $r=0.82$), one of the most commonly measured clinical outcomes for gait (Fig. 5C). Finally, the GDI showed a significant correlation with mJOA ($r=0.47$) and the lower extremity subscore of the mJOA ($r=0.48$) (Fig. 5E, Table 2).
Gait Measurement Repeatability
We analyzed the >>13,500 steps across 222 individuals recorded with the PBL system, and found high repeatability of 0.84 for the GDI of a single step (ICC2) and higher repeatability when averaging multiple steps together: 0.96 (ICC2k) (Table 2). ICCs for spatiotemporal metrics (cadence, DST) were lower-likely due to timing variability—but remained highly repeatable when averaged (ICC2k).
Responsiveness of Gait Measurement to Clinical Intervention
Smartphone-derived gait metrics (GDI and DST) demonstrated significant postoperative changes in individuals with cervical myelopathy (Wilcoxon signed-rank test, p<0.05), whereas patient-reported clinical scales (mJOA and mJOA-LE) did not show significant change (Table 2). Wilcoxon effect sizes were 0.24 for mJOA, 0.22 for mJOA-LE, -0.52 for DST, and 0.44 for GDI. In addition, paired SPM1D t-tests revealed a significant postoperative increase in knee flexion during initial swing (64–68 % of the gait cycle).
Discussion
@@
We present a clinically validated method for extracting biomechanically accurate kinematics from smartphone video in clinical settings. Across diverse populations, our single-camera system showed strong agreement with ground truth (typically $3^{\circ}$ Median Absolute Error). Gait quality metrics derived from these videos detected clinically relevant group differences and were responsive to clinical intervention. These findings suggest that a clinician, therapist, or medical assistant could feasibly perform gait analysis in under a minute to obtain an objective measurement of a patient’s gait impairment to track their recovery.
@@
Joint angle errors were generally low across all populations and settings (Table 1). We attribute the slightly higher Median Absolute Error in clinical populations (0.46 - 0.81 degrees, Fig. 4B, Fig. S1B, Table S2) to assistive devices such as canes, walkers, ankle foot orthoses, or gait belts (Fig. 3) and the fact that keypoint detection does not perform as well on prosthetic and orthotic devices [cimorelli_validation_2024]. Studies using OpenCap also report increases in joint angle errors in clinical populations [wang_evaluation_2025]. Intuitively, sagittal-plane angles were most accurate when recorded from sagittal viewpoints, and frontal-plane angles were most accurate from frontal viewpoints (Fig. 4C), giving clinicians a simple and actionable strategy for mitigating joint-angle error during data collection.
Our only dataset comparing PBL to Optical Motion Capture (OMC) is BML-MoVi, which only contains able-bodied controls. To further validate the biomechanical generalization of PBL, we also included a dataset with diverse etiologies of gait impairments assessed with our previously validated MMMC system [cotton_markerless_2023, cotton_differentiable_2025]. Although MMMC differs from OMC, multiple independent studies have demonstrated strong agreement between the two approaches [kanko_concurrent_2021, kanko_assessment_2021, riazati_absolute_2022, song_markerless_2023, wren_comparison_2023, horsak_repeatability_2024, dsouza_comparison_2024, unger_differentiable_2025]. Nonetheless, neither method constitutes a true biomechanical gold standard relative to biplanar fluoroscopy, due to known limitations including marker placement variability [uchida_conclusion_2022] and soft-tissue artifact [akbarshahi_non-invasive_2010]. The consistently strong correspondence observed between PBL and both OMC and MMMC across a diverse participant population supports the biomechanical accuracy and generalizability of the proposed approach (Table 1). While monocular video-based analysis is expected to exhibit some loss of accuracy relative to multiview systems, it substantially improves accessibility in clinical environments. Thus, the latter part of this study evaluates whether PBL collected in clinical settings is sufficiently accurate to demonstrate construct validity against established clinical metrics.
To demonstrate this, we showed that movement quality scores from PBL: (1) change with expected population changes, (2) correlate with clinical scores, and (3) are responsive to clinical intervention. First, GDI from PBL was decreased for inpatients with fall risk assessed by the BBS and further decreased for inpatients classified as high fall risk (Fig. 5B). Similarly, GDI was decreased for Transtibial LLPUs and further decreased for Transfemoral LLPUs, an intuitive result. Second, GDI correlated highly with the 10MWT (Fig. 5C), a widely used clinical outcome measure. The GDI also had a moderate correlation with the mJOA (Fig 5E, Table 2), which is expected considering the GDI is a continuous quantitative measure and the mJOA is a patient-reported ordinal scale. Regarding responsiveness, SPM1D analysis of kinematic traces showed significant changes in knee flexion following surgical intervention, establishing that our system can be used to analyze specific changes in movement and quantitatively track recovery. Quantitative metrics (DST and GDI) extracted from our smartphone-based system showed significant postoperative changes, whereas the mJOA did not (Table 2). This may suggest that smartphone-derived kinematic measures capture postoperative kinematics changes not reflected in patient-reported scores, but it is important to highlight these are measuring fundamentally different constructs. Thus, quantitative gait metrics (e.g., GDI, DST) complement patient-reported outcomes such as the mJOA, and we defer to future studies to investigate whether they may precede and causally influence patient outcomes [cotton_causal_2024].
@@
We focused on the GDI because it is a widely used summary of gait kinematics. However, we do not claim it is the optimal metric for every application. Instead, we anticipate that scalable gait analysis will enable new research to identify the most clinically meaningful measures. For example, we recently described a method for inferring torques and ground reaction forces from kinematics and saw that the ground reaction forces are sensitive to clinical history [cotton_kintwin_2025].
@@
The minimally clinically important difference for kinematic features of gait has not yet been established for many clinical conditions. While our accuracy is generally quite high, greater accuracy may be required for certain indications. For example, the 5 degrees of error we typically see at the ankle might be insufficient to detect subtle changes in plantarflexion spasticity or toe clearance from a single camera.
We observe heavy-tailed error distributions, with some trials exceeding 10° of error (Fig. S1C). Although occasional large errors and failure cases warrant further investigation, these errors occur infrequently and do not introduce systematic bias. Aggregated kinematic trajectories and summary metrics remain strongly correlated with clinically relevant measures (Table 2, Fig. 5), supporting the method’s reliability for population-level analysis and leaving subject-level clinical decision-making to future work. We tested whether fitting residuals—2D reprojection error or 3D keypoint error—could serve as reliability metrics. Their correlations with joint-angle error were weak (0.35 and 0.24, respectively; Fig. S5), underscoring the need for better confidence estimates in markerless motion capture. We have previously developed calibrated confidence intervals for monocular pose estimation without biomechanics [pierzchlewicz_platypose_2024] and for multi-view biomechanical estimation [cotton_biomechanical_2025]; extending these ideas to monocular biomechanics is an important direction for future work.
Other approaches have proposed deployable systems for human motion capture. Most well-known is OpenCap [uhlrich_opencap_2023], which uses two or three smartphones on static tripods to triangulate joint locations from video and fit biomechanics using the widely used OpenSim framework [delp_opensim_2007, werling_addbiomechanics_2023]. OpenCap reports a Mean Absolute Error of 4.1° during walking in 10 controls using two static cameras. In our control cohort (n=29), PBL achieves a comparable Mean Absolute Error of 4.32 (2.06)° using a single handheld moving camera. While this error was computed on similar populations and activities, we note this is not a direct comparison and leave that to future work. While more cameras will generally increase accuracy [cotton_biomechanical_2025], our results suggest that the difference between OpenCap with 2 camera on tripods and a single handheld smartphone is now small. Apart from accuracy, OpenCap and our PBL have different advantages: OpenCap provides a very user friendly platform, uses the widely accepted OpenSim package [delp_opensim_2007], and calculates joint kinetics. Our PBL approach presented here is markedly easier to deploy in clinical settings such as narrow hallways, small exam rooms, and busy gym spaces. The same simplicity also enables potential use in patients’ homes with recordings performed by caregivers. Extending our method to include joint kinetics is feasible with future work such as [cotton_kintwin_2025]. While our pipeline is not real-time, per-trial runtimes are compatible with same-day or visit-level clinical analysis.
Two recent approaches estimate biomechanics from monocular video: BioPose [koleini_biopose_2025] and Human Skeleton and Mesh Recovery (HSMR) [xia_reconstructing_2025]. Both regress pose directly from images, making them faster than our pipeline. However, neither was trained or evaluated on clinical populations, so their performance on clinical gait—often far outside the distribution of healthy movement (Fig. 5A)—is uncertain. Our method instead uses trajectory optimization with in-loop biomechanical model scaling applied consistently across trials, aligning more closely with established biomechanical reconstruction practices. It can also integrate wearable sensors [peiffer_fusing_2024] to address occlusion and high-speed motion. For healthy controls, our kinematic accuracy is similar to BioPose (<3.9° Mean Absolute Error on MoVi dataset), though the comparison is incomplete because BioPose does not specify which joints contribute to its reported errors. HSMR does not report joint-angle errors. Ultimately, broader development of AI-driven biomechanics tools will expand clinical movement analysis, and large resulting datasets will support data-driven, precision rehabilitation strategies [cotton_causal_2024].
@@
In conclusion, we present a method for extracting accurate and biomechanically interpretable joint kinematics from smartphone videos. We validate this method in inpatient, outpatient therapy, and clinical settings, finding it to be highly usable and robust—even in the presence of nearby therapists and assistive devices. Importantly, gait scores extracted from this approach correlate well with commonly used clinical scales such as the 10MWT, BBS, and mJOA, and respond to clinical intervention. This approach enables increased fidelity in clinical and at-home monitoring of movement impairments, paving the way for a big data revolution in movement science.
@@ Acknowledgements.
OMC Dataset: BML-MoVi
The publicly available BML-MoVi [ghorbani_movi_2021] dataset contains synchronous optical motion capture (OMC) data and smartphone RGB video data of 90 able-bodied participants performing a variety of everyday and sports movements. The OMC data contains a dense set of 87 markers, suitable for reconstructing biomechanics recorded at 120 Hz.
For the OMC data, we fit our model using only the 3D keypoint loss as the markers had already been triangulated and cleaned, following the approach from [cotton_differentiable_2025]. For the smartphone data, we apply the 3D and 2D keypoint loss following the method we use on the PBL data and describe below.
@@
This study was approved by the Northwestern University Institutional Review Board. Participants provided informed consent for their videos to be stored, processed, and de-identified videos to be shared in publication. Data were collected in our laboratory using handheld smartphone video [cimorelli_validation_2024], an instrumented GaitRite walkway, and an MMMC system [cotton_markerless_2023]. The dataset includes recordings from 33 control participants, 48 lower limb prosthesis users (LLPUs), and 73 individuals with a history of neurological injury, including 6 pediatric cases. All recording sessions consisted of overground walking along the GaitRite walkway (8 m) while a researcher followed the participant with smartphone.
@@
Portable Biomechanics Laboratory (PBL) All smartphone RGB video was captured on a Samsung Galaxy S8 or Samsung Galaxy S20 Ultra at 30 Hz and 1080 $\times$ 1920 resolution [cimorelli_validation_2024] using a custom application. This application simultaneously logs smartphone RGB video, and if available, depth video, internal phone sensor logs (gyroscope, magnetometer, accelerometer, and fused orientation estimate), and wearable sensor data [cotton_wearable_2019]. During walking trials, a researcher held the phone in portrait orientation mounted on a 3-axis gimbal and followed the participant from behind at a mostly sagittal, but partially oblique view. The gimbal was used solely to improve visualizations; the method remains accurate without it, though the resulting video may be slightly shakier. FSST was collected at a frontal view in front of the participant, and TUG was collected from a sagittal view with the camera rotating to follow the participant.
The video and sensor data from the smartphone application was uploaded to a secure cloud server for storage and later downloaded into PosePipe [cotton_posepipe_2022], an open source application implementing a variety of computer vision tasks such as person detection, annotation, and 2D and 3D keypoint detection. PosePipe tracks all data and analyses in a MySQL database using DataJoint [yatsenko_datajoint_2015]. This makes it feasible to manage the 100s of thousands of videos we have collected.
Multi-View Markerless Motion Capture (MMMC) Recordings in the laboratory dataset employed the Multi-View Markerless Motion Capture (MMMC) system which consists of 8-12 FLIR BlackFly S GigE cameras which acquire synchronized RGB video at 29 fps. Cameras were arranged such that at least three cameras covered participants at all points during a recording. A detailed description of this system can be found in [cotton_markerless_2023]. For this system, we reconstructed biomechanical fits using methods described in [cotton_differentiable_2025]. In total, the dataset includes 1,267,586 frames across 2,206 trials from 143 participants—comprising 1,056,861 frames from 2,028 trials of 116 participants recorded with both the PBL and MMMC systems, and an additional 210,725 frames from 178 trials of 27 participants recorded with a single MMMC camera—amounting to 11.7 hours of video.
@@
This study was approved by the Northwestern University and Washington University in St. Louis Institutional Review Boards. Participants provided informed consent for their videos to be stored, processed, and de-identified videos to be shared in publication. To validate our methods in a real-world clinical environment, we deployed our PBL system [cimorelli_validation_2024] in two outpatient clinics. The first was a neurosurgery clinic treating patients with cervical myelopathy (CM)—a degenerative condition of the cervical spine that often impairs gait and balance. During routine pre- and post-operative visits for spinal decompression, participants were asked to walk three times at a self-selected pace and three times at a fast pace down the clinic hallway while being recorded from behind by a clinician. At each visit, patients completed the modified Japanese Orthopedic Association (mJOA) questionnaire [yonenobu_interobserver_2001, kato_comparison_2015] to assess their symptoms. The CM cohort includes 911 videos from 55 patients collected at multiple timepoints: pre-surgery, 6 weeks post-surgery, 3 months post-surgery, and 1 year post-surgery. 17 of these participants have sessions before and after surgery, which were used in our sensitivity analysis.
@@
We employed MeTRAbs-ACAE [sarandi_learning_2023] to detect 2D and 3D virtual marker locations from RGB videos. As recommended by the author, we approximated confidence by measuring the standard deviation of each 3D joint location estimated from 10 different augmented versions of each video frame. This was converted to a confidence estimate using a sigmoid function with a half maximum at 30 mm and a width of 10 mm. This results in a confidence score $c(t,j)\in[0,1]$ for each timepoint $t$ and joint $j$. For training, we only used frames where the person was fully in the video view, allowing for short periods (<<1.0) seconds of partial coverage, reducing total dataset size to 12.4 hours. For a 6 second video with one person in frame, MeTRAbs-ACAE took 10 seconds to run on a NVIDIA A100 GPU and 1 minute to run on a NVIDIA GeForce RTX 3050 Laptop GPU. This processing time will scale linearly with video time.
@@
Smartphone and MMMC reconstructions utilized a biomechanically-grounded model implemented in Mujoco from LocoMujoco [al-hafez_locomujoco_2023], originally based on an OpenSim model [hamner_muscle_2010]. Mujoco’s GPU acceleration engine, MJX, supports parallelizing forward kinematic passes of this model. We have previously optimized site locations for the 87 MoVi keypoints on a similar model [cotton_optimizing_2023]. We also removed collisions other than the feet, added a neck joint with 3 degrees of freedom, and extended the hip flexion and extension ranges. Model scaling is controlled by 8 scaling parameters: overall size, the pelvis, left thigh, left leg and foot, right thigh, right leg and foot, the left arm, and the left leg. We follow notation used by the skinned multi-person linear model (SMPL) [loper_smpl_2015] and represent our GPU-accelerated, biomechanically-ground, forward kinematic equation as:
@@
We represent the kinematic trajectory for each trial as a learned implicit function, $f_{\theta}$, implemented as a multi-layer perceptron that takes time as an input using sinusoidal positional encoding [vaswani_attention_2017] and outputs joint angles $\theta(t)$. In addition to the 40 joint angles output from the final layer, we also output 3 parameters representing the orientation of the smartphone in global space $\mathbf{r}$.